Question 1. Consider 3 gene loci in the pea plant. Locus 1 affects seed color; allele G for green, g for yellow. Locus 2 affects seed shape; allele R for round, r for wrinkled. Locus 3 affects flower color; allele B for red, b for white. Each uppercase allele is completely dominant to their corresponding lowercase, recessive allele. Each locus is on a different non-homologous chromosome; so, the loci segregate or sort independently from each other during meiosis. Two parent pea plants, each of genotype GgRrBb, are mated with each other. That mating produced 1000 progeny.

A) What fraction of the gametes of the parents are expected to contain only recessive alleles? Type your answer as a fraction in simplest form (e.g.,1/2)

B) How many phenotypic classes are expected among the progeny? Type your answer as an integer (e.g.,6)

C) What fraction of the progeny are expected to have the combination green seed, round seed, and red flowers? Type your answer as a fraction in simplest form (e.g.,1/2)

D) How many progeny are expected to have a combination of yellow seed, wrinkled seed, and white flowers? Give your answer rounded to the nearest whole number (e.g.,6)

E) How many genotypic classes are expected among the progeny?​ Type your answer as an integer (e.g.,6)​​​​​​
Answer #1

Answer:

A). 1/8

Explanation:

GgRrBb produce 8 types of gametes with equal probability, i.e., 1/8.

GRB=1/8

GRb = 1/8

GrB = 1/8

Grb = 1/8

gRB=1/8

gRb = 1/8

grB = 1/8

grb = 1/8

B). 8 types of phenotypes will be produced.

Explanation:

Number phenotypes = 2^n = 2^3 = 8

n = Number of genes

C).

The fraction of the progeny are expected to have the combination green seed, round seed, and red flowers = G_R_B_ = ¾ * ¾ * ¾ = 27/64

Explanation:

GgRrBb x GgRrBb ---Parents

Gg X Gg = GG (1/4), Gg (1/2) & gg (1/4)

Rr x Rr = RR (1/4), Rr (1/2) & rr (1/4)

Bb x Bb = BB (1/4), Bb (1/2) & bb (1/4)

Number of progeny are expected to have a combination of yellow seed, wrinkled seed, and white flowers = gg rr bb = ¼ * ¼ * ¼ * 1000 = 16

E). 27

Genotypic ratio: 1:2:2:2:4:8:4:4:2:2:1:4:1:2:4:2:1:2:1:4:2:2:1:2:1:2:1.